OpenStudy (1230kent):

Jane , Andre and Maria picked 840 pounds of apples. Andre picked 3 times as many as Maria. Jane picks 2 times as many as Andre. How many pounds did Andre pick.

OpenStudy (anonymous):

Make some equations.. jane=j andre=a maria=m j+a+m=840 a=3m j=2a any idea what to do now?

OpenStudy (anonymous):

Let the amount Maria picks = m. Andre picks 3 times as many as Maria so Andre = 3m Jane picks 2 times as many as Andre so Jane = 2(3m) = 6m Therefore; m + 3m + 6m = 840 total pounds. Rearrange to find m; 10m = 840 840/10 = 84. So Maria picks 84 pounds, Andre picks 252 pounds and Jane picks 504 pounds.

OpenStudy (1230kent):

I am in 5th grade trying to learn this and that is not how my teacher showed us. Where did you get 3m from oh wait I see it now. Can you walk me through it I get it kinda. thank you for your help

OpenStudy (anonymous):

M is just a variable for Maria. What do you need help with exactly? O:

OpenStudy (1230kent):

we use boxes because we are just learning this

OpenStudy (anonymous):

Boxes? O: Hmm I don't think we ever learnt that way so not sure I can help o-o But to clarify when I wrote '3m' it means 3*m (3 times m) and 2(3m) means 2*3*m (2 times 3 times m)